OpenStudy (anonymous):

hat is the slope of a line that passes through the points (9, 81) and (6, 36)?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

A line on a coordinate grid passes through the points (-2, 5) and (-3, 6). What is the slope of the line

OpenStudy (anonymous):

-1

OpenStudy (igreen):

To find the slope of two points plug it into the slope formula: \(m = \dfrac{y_2-y_1}{x_2-x_1}\) In this case: \(y_2 = 36\) \(y_1 = 81\) \(x_2 = 6\) \(x_1 = 9\) So: \(m = \dfrac{36-81}{6-9}\)

OpenStudy (igreen):

@tmoneycash4 You're not supposed to give direct answers.

OpenStudy (anonymous):

sorry you do really need to work the promblems out on paper before hand thats what i did

OpenStudy (anonymous):

k

OpenStudy (anonymous):

can you help me on my problem @iGreen

OpenStudy (anonymous):

@phi i neeed hlp plzz

OpenStudy (phi):

what is the question ?

OpenStudy (anonymous):

messges meh plz

OpenStudy (anonymous):

and ill puti in yo inbox

OpenStudy (phi):

Please make a new post.

OpenStudy (anonymous):

ok